# Rafael A. Berri1 Giancarlo Lucca2    Fabian C. Cardoso3

### v1

@@ I Introduction
In the modern world, computers are extremely important and are used in nearly every aspect of daily life. They are continuously advancing, with constant improvements in their performance and capabilities, making their efficient use increasingly necessary and important [NRC2011].
In particular, the CPU (Central Processing Unit) is evolving at a rapid pace, with annual improvements in processing speed and memory storage, known in the CPU context as cache. The cache has different levels of speed and size, whereas other components have not evolved at the same rate [nyberg2021investigating]. Storage speed outside the CPU is significantly slower, and in this paper, we will focus on RAM, which is approximately 100 times slower than the CPU [carvalho2002gap].
@@
OOP is not efficient in terms of cache usage, as it has a greater impact on performance and scalability, whereas DOD does not exhibit these issues [nikolov2018oop]. OOP consists of using object classes for data manipulation and well-encapsulated functions with multiple abstraction layers [stroustrup1988object]. In contrast, DOD makes better use of data by separating it from the code, thus improving data access and allowing for greater scalability and easier code maintenance [joshi2007data].
@@ II-A Articles and Books
Computer architecture is a vast and constantly evolving field, with numerous significant contributions over the decades. In this context, the book ”Computer Architecture” by Hennessy and Patterson [patterson1988] highlights the importance of memory in computing, particularly the role of cache in enhancing overall system performance. They argue that systems with more efficient caches tend to perform better overall, even though they do not provide a study specifically focused on cache, supporting their assertion through a broad analysis of computer architectures.
The relevance of cache is corroborated by various other studies and publications. For example, Culler et al. [culler1998], in ”Parallel Computer Architecture: A Hardware/Software Approach,” emphasize that the performance of parallel systems is crucial to avoiding bottlenecks and demonstrate how effective cache utilization can significantly improve system performance. Similarly, Drepper [drepper2007], in ”What Every Programmer Should Know About Memory,” discusses how efficient cache management can reduce memory access time, thereby improving program performance.
Leslie Lamport, in his seminal paper ”Time, Clocks, and the Ordering of Events in a Distributed System” [lamport1978], introduces concepts of continuity and synchronization in distributed systems, emphasizing the importance of cache in reducing latency and enhancing performance. This perspective is complemented by Gene Amdahl [amdahl1967], who mentions the significance of cache in minimizing memory access latency in his work on Amdahl’s Law, although his primary focus is on parallelism.
The impact of cache on performance is also addressed by Patterson et al. [patterson1988] in ”A Case for Redundant Arrays of Inexpensive Disks (RAID),” where they discuss how efficient caches can improve the performance of data storage systems. Dean and Ghemawat [dean2004], in ”MapReduce: Simplified Data Processing on Large Clusters,” highlight the importance of intelligent cache usage to enhance the performance of processing nodes in clusters, which is essential for the parallel processing of large datasets.
In ”High-Performance Concurrency Control Mechanisms for Main-Memory Databases,” Larson et al. [larson2011] analyze the use of multi-threading to increase cache efficiency in highly concurrent database environments, demonstrating that an efficient cache is crucial for reducing latency and increasing throughput. Similarly, Bienia et al. [bienia2008], in ”The PARSEC Benchmark Suite: Characterization and Architectural Implications,” present benchmarks that show how the structure and effectiveness of the cache are fundamental to achieving good performance in parallel applications dealing with large volumes of data.
@@
One of the most vocal critics of OOP is Joe Armstrong, the creator of the Erlang programming language. In his article ”Why OO Sucks” [armstrong2003], Armstrong argues that the strong interdependence between objects and the complexity of class hierarchies can lead to rigid systems that are difficult to modify. He asserts that DOD, by focusing on data structure and flow, offers greater flexibility and simplicity. This approach can result in systems that are easier to maintain and scale, as the programmer’s logic becomes clearer and more direct.
Paul Graham, in ”Hackers & Painters” [graham2004], also criticizes OOP for the performance overhead it can cause. He argues that abstraction and encapsulation, fundamental to OOP, can perform worse compared to more direct approaches such as procedural programming or DOD. The simplicity of DOD can lead to more efficient and easier-to-understand code, avoiding the excessive complexity that often accompanies OOP.
Robert C. Martin, in ”Clean Code” [martin2009], acknowledges that OOP can create unnecessary complexity if not used cautiously. While he does not completely discard OOP, Martin suggests that maintaining simplicity and avoiding complex class hierarchies are essential practices. DOD, by its nature, tends to avoid these problems by focusing on data structures in a more linear and direct manner.
@@ II-B Related Work
The article by Pellacini et al. [pellacini2019yocto], ”Yocto/GL: A Data-Oriented Library For Physically-Based Graphics,” emphasizes the use of Data-Oriented Design (DOD) as a way to achieve simplicity, composability, and scalability in computer graphics. The authors contrast this approach with Object-Oriented Design (OOD), frequently used in other libraries, which they argue can lead to complexity and performance issues, especially when handling large datasets. Specifically, Yocto/GL uses free functions that operate on basic data types, value semantics for assignments, and index references instead of pointers. These design choices aim to improve code readability, facilitate the composition of different functionalities, and promote a more memory-efficient data layout.
@@
In ”Performance of Data-Oriented Design with Unity DOTS,” Venâncio [venancio2023desempenho] compares Data-Oriented Design (DOD) with Object-Oriented Design (OOD) in game development using the Unity engine and the DOTS package. The author argues that DOD, by focusing on processing and data, facilitates the implementation of multi-threading, resulting in better CPU utilization and consequently improved performance. The study includes a tank battle simulator, ”Metal Rain,” to test both approaches. The results show that DOD, with DOTS, achieves higher frames per second (FPS) and better CPU core utilization, especially in scenarios with many elements. It is important to note that the author observes that in cases of GPU bottlenecks, the benefits of DOD in terms of FPS may be less apparent. Although the study does not delve deeply into cache optimization, the author mentions that DOD generally organizes data in arrays, which favors cache memory usage.
@@
In ”Exploiting Cache Locality to Speedup Register Clustering,” Fontana et al. [fontana2017exploiting] argue that data organization directly influences algorithm performance, especially in tasks such as register clustering, which involve large volumes of data. The authors compare Data-Oriented Design (DOD) with traditional Object-Oriented Design (OOD), demonstrating that DOD, by organizing data into separate arrays per property, reduces cache space wastage and improves cache spatial locality. This optimization translates into fewer cache misses and consequently faster execution times in both sequential and parallel (multi-threaded) implementations. The article demonstrates, through experiments with the K-means algorithm, that the data organization proposed by DOD accelerates processing, particularly in high-memory-use scenarios, making it a promising approach for Electronic Design Automation (EDA) tools, which are software tools used to design and produce electronic systems, from computer chips to printed circuit boards.
@@
In ”Investigating the Effect of Implementing Data-Oriented Design Principles on Performance and Cache Utilization,” Nyberg et al. [nyberg2021investigating] argue that Data-Oriented Design (DOD) principles can significantly enhance performance in game development environments, primarily through optimized CPU cache utilization. The author explores the removal of runtime polymorphism, iteration over contiguous data, and reduction of data size as core DOD principles that directly influence cache performance. Through experiments with a test program simulating movement and collision detection, Nyberg demonstrates that these DOD optimizations result in faster CPU times and lower cache miss rates compared to traditional object-oriented designs. The study highlights the importance of data locality and efficient cache usage for high-performance games, especially in data-intensive scenarios. However, the author acknowledges that the scope of the research is limited to single-threaded tests and suggests further investigation in multi-threaded environments as a topic for future research.
@@
Wingqvist et al. [wingqvist2022], in ”Evaluating the Performance of Object-Oriented and Data-Oriented Design with Multi-Threading in Game Development,” developed a simple game using two programming design patterns, Object-Oriented Design (OOD) and Data-Oriented Design (DOD), to compare their performance in terms of runtime and cache utilization, both in single-threaded and multi-threaded scenarios. They ran the game on different systems with various numbers of cores and cache sizes and found that DOD consistently outperformed OOD across all configurations, exhibiting significantly faster execution times (achieving a speedup of up to 13.25 times) and lower cache miss rates (with an improvement of up to 5.57 times in multi-threaded scenarios). The authors concluded that DOD is particularly beneficial for game development, especially on modern multi-core systems, as it utilizes CPU resources more efficiently. However, they acknowledged that DOD is more complex to implement than OOD and suggested future research with large-scale games to further validate their findings.

### v2

# Gabriel M. Arantes1, Giancarlo Lucca2, Eduardo N. Borges1, Richard F. Pinto1, Bruno L. Dalmazo1 and Rafael A. Berri1

@@ I Introduction
In the modern world, computers are extremely important and are used in nearly every aspect of daily life. They are continuously advancing, with constant improvements in their performance and capabilities, making their efficient use increasingly necessary and important [5].
In particular, the CPU (Central Processing Unit) is evolving at a rapid pace, with annual improvements in processing speed and memory storage, known in the CPU context as cache. The cache has different levels of speed and size, whereas other components have not evolved at the same rate [16]. Storage speed outside the CPU is significantly slower, and in this paper, we will focus on RAM, which is approximately 100 times slower than the CPU [4].
@@
OOP is not efficient in terms of cache usage, as it has a greater impact on performance and scalability, whereas DOD does not exhibit these issues [15]. OOP consists of using object classes for data manipulation and well-encapsulated functions with multiple abstraction layers [19]. In contrast, DOD makes better use of data by separating it from the code, thus improving data access and allowing for greater scalability and easier code maintenance [11].
@@ II-A Articles and Books
Computer architecture is a vast and constantly evolving field, with numerous significant contributions over the decades. In this context, the book ”Computer Architecture” by Hennessy and Patterson [17] highlights the importance of memory in computing, particularly the role of cache in enhancing overall system performance. They argue that systems with more efficient caches tend to perform better overall, even though they do not provide a study specifically focused on cache, supporting their assertion through a broad analysis of computer architectures.
The relevance of cache is corroborated by various other studies and publications. For example, Culler et al. [6], in ”Parallel Computer Architecture: A Hardware/Software Approach,” emphasize that the performance of parallel systems is crucial to avoiding bottlenecks and demonstrate how effective cache utilization can significantly improve system performance. Similarly, Drepper [8], in ”What Every Programmer Should Know About Memory,” discusses how efficient cache management can reduce memory access time, thereby improving program performance.
Leslie Lamport, in his seminal paper ”Time, Clocks, and the Ordering of Events in a Distributed System” [12], introduces concepts of continuity and synchronization in distributed systems, emphasizing the importance of cache in reducing latency and enhancing performance. This perspective is complemented by Gene Amdahl [1], who mentions the significance of cache in minimizing memory access latency in his work on Amdahl’s Law, although his primary focus is on parallelism.
The impact of cache on performance is also addressed by Patterson et al. [17] in ”A Case for Redundant Arrays of Inexpensive Disks (RAID),” where they discuss how efficient caches can improve the performance of data storage systems. Dean and Ghemawat [7], in ”MapReduce: Simplified Data Processing on Large Clusters,” highlight the importance of intelligent cache usage to enhance the performance of processing nodes in clusters, which is essential for the parallel processing of large datasets.
In ”High-Performance Concurrency Control Mechanisms for Main-Memory Databases,” Larson et al. [13] analyze the use of multi-threading to increase cache efficiency in highly concurrent database environments, demonstrating that an efficient cache is crucial for reducing latency and increasing throughput. Similarly, Bienia et al. [3], in ”The PARSEC Benchmark Suite: Characterization and Architectural Implications,” present benchmarks that show how the structure and effectiveness of the cache are fundamental to achieving good performance in parallel applications dealing with large volumes of data.
@@
One of the most vocal critics of OOP is Joe Armstrong, the creator of the Erlang programming language. In his article ”Why OO Sucks” [2], Armstrong argues that the strong interdependence between objects and the complexity of class hierarchies can lead to rigid systems that are difficult to modify. He asserts that DOD, by focusing on data structure and flow, offers greater flexibility and simplicity. This approach can result in systems that are easier to maintain and scale, as the programmer’s logic becomes clearer and more direct.
Paul Graham, in ”Hackers & Painters” [10], also criticizes OOP for the performance overhead it can cause. He argues that abstraction and encapsulation, fundamental to OOP, can perform worse compared to more direct approaches such as procedural programming or DOD. The simplicity of DOD can lead to more efficient and easier-to-understand code, avoiding the excessive complexity that often accompanies OOP.
Robert C. Martin, in ”Clean Code” [14], acknowledges that OOP can create unnecessary complexity if not used cautiously. While he does not completely discard OOP, Martin suggests that maintaining simplicity and avoiding complex class hierarchies are essential practices. DOD, by its nature, tends to avoid these problems by focusing on data structures in a more linear and direct manner.
@@ II-B Related Work
The article by Pellacini et al. [18], ”Yocto/GL: A Data-Oriented Library For Physically-Based Graphics,” emphasizes the use of Data-Oriented Design (DOD) as a way to achieve simplicity, composability, and scalability in computer graphics. The authors contrast this approach with Object-Oriented Design (OOD), frequently used in other libraries, which they argue can lead to complexity and performance issues, especially when handling large datasets. Specifically, Yocto/GL uses free functions that operate on basic data types, value semantics for assignments, and index references instead of pointers. These design choices aim to improve code readability, facilitate the composition of different functionalities, and promote a more memory-efficient data layout.
@@
In ”Performance of Data-Oriented Design with Unity DOTS,” Venâncio [20] compares Data-Oriented Design (DOD) with Object-Oriented Design (OOD) in game development using the Unity engine and the DOTS package. The author argues that DOD, by focusing on processing and data, facilitates the implementation of multi-threading, resulting in better CPU utilization and consequently improved performance. The study includes a tank battle simulator, ”Metal Rain,” to test both approaches. The results show that DOD, with DOTS, achieves higher frames per second (FPS) and better CPU core utilization, especially in scenarios with many elements. It is important to note that the author observes that in cases of GPU bottlenecks, the benefits of DOD in terms of FPS may be less apparent. Although the study does not delve deeply into cache optimization, the author mentions that DOD generally organizes data in arrays, which favors cache memory usage.
@@
In ”Exploiting Cache Locality to Speedup Register Clustering,” Fontana et al. [9] argue that data organization directly influences algorithm performance, especially in tasks such as register clustering, which involve large volumes of data. The authors compare Data-Oriented Design (DOD) with traditional Object-Oriented Design (OOD), demonstrating that DOD, by organizing data into separate arrays per property, reduces cache space wastage and improves cache spatial locality. This optimization translates into fewer cache misses and consequently faster execution times in both sequential and parallel (multi-threaded) implementations. The article demonstrates, through experiments with the K-means algorithm, that the data organization proposed by DOD accelerates processing, particularly in high-memory-use scenarios, making it a promising approach for Electronic Design Automation (EDA) tools, which are software tools used to design and produce electronic systems, from computer chips to printed circuit boards.
@@
In ”Investigating the Effect of Implementing Data-Oriented Design Principles on Performance and Cache Utilization,” Nyberg et al. [16] argue that Data-Oriented Design (DOD) principles can significantly enhance performance in game development environments, primarily through optimized CPU cache utilization. The author explores the removal of runtime polymorphism, iteration over contiguous data, and reduction of data size as core DOD principles that directly influence cache performance. Through experiments with a test program simulating movement and collision detection, Nyberg demonstrates that these DOD optimizations result in faster CPU times and lower cache miss rates compared to traditional object-oriented designs. The study highlights the importance of data locality and efficient cache usage for high-performance games, especially in data-intensive scenarios. However, the author acknowledges that the scope of the research is limited to single-threaded tests and suggests further investigation in multi-threaded environments as a topic for future research.
@@
Wingqvist et al. [21], in ”Evaluating the Performance of Object-Oriented and Data-Oriented Design with Multi-Threading in Game Development,” developed a simple game using two programming design patterns, Object-Oriented Design (OOD) and Data-Oriented Design (DOD), to compare their performance in terms of runtime and cache utilization, both in single-threaded and multi-threaded scenarios. They ran the game on different systems with various numbers of cores and cache sizes and found that DOD consistently outperformed OOD across all configurations, exhibiting significantly faster execution times (achieving a speedup of up to 13.25 times) and lower cache miss rates (with an improvement of up to 5.57 times in multi-threaded scenarios). The authors concluded that DOD is particularly beneficial for game development, especially on modern multi-core systems, as it utilizes CPU resources more efficiently. However, they acknowledged that DOD is more complex to implement than OOD and suggested future research with large-scale games to further validate their findings.